Several images are blank, likely with the dome completely in the way
The centering of the field seems to drift from night to night
At high hour angle, there is significant trailing in the images. There is still notable “coma” in images near zenith
Several have noticed the following behavior in the last few nights with the RCT:
I believe these are all symptoms of the encoder slippage and worm gear oscillation. I believe they needing cleaning/re-greasing. I’ve put in a service request with KPNO. I still think part of the PSF is mirror support related, but we wont know for sure until we eliminate the other sources of confusion.
Tonight I was able to zero point the telescope, and test the guider. It appears that the pointing tolerance for accurate guiding is about 15″, or ~20 pix from the absolute center of the chip. Guiding works better now, but it does not improve the PSF substantially.
Data archive is back. There may be a little lag in getting data from the last few nights.
They are installing new power outlets in our server room. The computers, including mnem, will be down until 10/9.
The mirror support compressor’s pressure switch, which activates and deactivates the pump, burned out. It was replaced with a new switch, one with a bit more control on its activation pressure, which should reduce the wear on this part.
The problem with the mirror support compressor seems to be resolved, and the telescope should be working normally. I will check the focus and pointing again tonight, then resume the scheduler.